### v1

# Uncovering the Functional Roles of Nonlinearity in Memory

Abstract
Memory and long-range temporal processing are core requirements for sequence modeling tasks across natural language processing, time-series forecasting, speech recognition, and control. While nonlinear recurrence has long been viewed as essential for enabling such mechanisms, recent work suggests that linear dynamics may often suffice. In this study, we go beyond performance comparisons to systematically dissect the functional role of nonlinearity in recurrent networks—identifying both when it is computationally necessary, and what mechanisms it enables. We use Almost Linear Recurrent Neural Networks (AL-RNNs), which allow fine-grained control over nonlinearity, as both a flexible modeling tool and a probe into the internal mechanisms of memory. Across a range of classic sequence modeling tasks and a real-world stimulus selection task, we find that minimal nonlinearity is not only sufficient but often optimal, yielding models that are simpler, more robust, and more interpretable than their fully nonlinear or linear counterparts. Our results provide a principled framework for selectively introducing nonlinearity, bridging dynamical systems theory with the functional demands of long-range memory and structured computation in recurrent neural networks, with implications for both artificial and biological neural systems.
1 Introduction
Modeling sequences with temporal dependencies remains a core challenge in machine learning, vital for domains such as language, speech, control, and time-series forecasting [28, 65, 35, 36]. These tasks often depend on long-range memory—requiring models to store, transform, and retrieve information across time. Classical results have long pointed to the necessity of nonlinear recurrence for such operations, showing that recurrent neural networks (RNNs) with nonlinear dynamics are both universal approximators and Turing complete [59].
Yet, recent work has revived interest in linear recurrent models, particularly due to their computational advantages. Linear systems support parallel computation and scale efficiently [43], and when equipped with expressive input and output mappings, they have been shown to perform surprisingly well on long-horizon tasks [24, 48]. These findings revive a deeper question: To what extent is nonlinearity necessary for temporal computation, and what role does it actually play?
This question is especially relevant when considering interpretability. Linear systems, despite their dynamical limitations [61], are analytically tractable and structurally transparent, making them appealing in settings where understanding the internal dynamics of a model is as important as its raw performance (as, for instance, in neuroscience [62, 16]). Nonlinear systems, by contrast, afford richer dynamics, but typically introduce a higher degree of complexity.
Reflecting this trade-off, the landscape of sequence models has grown into a diverse collection of architectures combining linearity and nonlinearity in different ways [50], for instance through mechanisms such as input-dependent parameter nonlinearities [22] or gating mechanisms [44]. These innovations, while empirically successful, often blur the functional roles of linear and nonlinear components, obscuring their individual contributions to sequence modeling. This lack of clarity motivates our investigation: we seek to disentangle how and when nonlinearity is functionally necessary, and to better understand its computational benefits across diverse tasks.
We tackle this question using the Almost-Linear RNN (AL-RNN), a model initially proposed for dynamical systems reconstruction (DSR) [9]. Here, we repurpose the AL-RNN as a natural tool to investigate a wide class of sequence modeling tasks. The AL-RNN is uniquely suited for this purpose due to its ability to adjust the degree of nonlinearity in its recurrence, allowing for a systematic exploration of where and when nonlinearity is functionally necessary. After inference, the AL-RNN offers dynamical and computational interpretability: First, its piecewise linear (PWL) structure allows analytical access to its dynamical objects such as fixed points, stability properties, and bifurcations [17]. Second, the AL-RNN naturally generates a symbolic representation of its internal state transitions, structured through a bitcode that reflects its nonlinear switching mechanisms [9].
In our study, we apply the AL-RNN across a diverse set of classic memory tasks, including neural recordings from a rodent performing a stimulus selection task, systematically exploring scenarios where linearity suffices, minimal nonlinearity is beneficial, or strong nonlinearity is required. A central finding of our study is that sparsely applied nonlinearity—introduced selectively within an otherwise linear system—can match or sometimes even exceed the performance of fully nonlinear models while retaining strong interpretability. In doing so, we reveal task-specific nonlinear mechanisms such as gating, rule-based integration, and memory-dependent transients, all embedded within a mostly linear dynamical backbone. Our study sheds light on the minimum necessary conditions for memory in recurrent models, offering a principled framework guiding the design of sequence models that are both computationally efficient and analytically tractable.
2 Related Work
The ability to represent, retain, and manipulate memory over time has been a central challenge in sequence modeling. Early RNNs, such as those proposed by Elman [18] and Jordan [30], introduced the paradigm of learning temporal dependencies through hidden state recurrence. However, these simple RNNs struggled to capture long-range dependencies due to vanishing and exploding gradients [6]. This limitation was addressed by architectures like Long Short-Term Memory (LSTM) networks [28] and Gated Recurrent Units [13]. Such gated RNNs, and more recent extensions such as Long Expressive Memory [56], established gated recurrence as a cornerstone of sequence modeling, supporting stable memory retention across extended time spans. These models successfully mitigated gradient issues, but the introduction of nonlinearity through gating also increased the complexity of their internal state dynamics. To address this, an alternative line of work has focused on combining the expressivity of nonlinear dynamics with the analytical tractability of linear models. This gave rise to piecewise linear and switching models—such as switching linear dynamical systems [19, 38, 37] and piecewise linear recurrent neural networks (PLRNNs) [15, 8, 9]. These architectures segment the state space into locally linear regions, enabling rich dynamical behaviors through structured transitions between subspaces while preserving mathematical tractability.
While the computational complexity of RNNs scales with $\mathcal{O}(T)$, $T$ being the sequence length, the inherently sequential nature of RNNs makes their training inefficient to parallelize on modern GPU hardware. The Transformer architecture [65] has therefore largely replaced recurrence with global self-attention across many sequence-modeling tasks. Self-attention enables parallel sequence processing without relying on a persistent hidden state. This parallelism, however, comes at the cost of computational efficiency, scaling with $\mathcal{O}(T^{2})$. Further, it models memory only within a finite context window. As a result, Transformers lack causal, evolving memory, making them harder to interpret and less inherently efficient for very long-range dependencies [50].
This quadratic scaling has motivated the exploration of more efficient alternatives for long-sequence modeling. Recent interest has focused on structured linear state-space models (SSMs). These models implement linear recurrence, which can be unrolled using spectral methods or convolutions. Advanced parallel scanning techniques, such as Blelloch’s scan [7], enable linear RNNs to achieve highly efficient parallelization, reducing complexity to $\mathcal{O}(\log T)$ [43]. Theoretical advancements have even demonstrated that deep linear RNNs with nonlinear mixing layers are universal approximators of regular sequence maps, broadening their applicability to complex temporal tasks [48, 49]. Modern SSMs like S4 [24], grounded in HiPPO theory [23], leverage these principles to match, or even surpass, Transformers on long-range sequence benchmarks [63]. Recent extensions have introduced more efficient initialization schemes and streamlined designs [25, 60, 26], enhancing both stability and performance. Some models, like Mamba [22] and Gated SSMs [44], reintroduce nonlinearity through structured gating mechanisms while preserving the parallelizable structure of SSMs.
3 Method
Almost-Linear Recurrent Neural Networks
To dissect the roles of linear and nonlinear recurrence, we adopt an SSM-inspired architecture: the recently proposed AL-RNN [9]. In this model, latent dynamics evolve under a combination of linear and PWL transition functions, modulated by external inputs. Nonlinearities are applied only to a subset $P\leq M$ of the hidden state, making AL-RNNs a simplified variant of piecewise linear RNNs [15, 8].
An AL-RNN is defined by:
$\bm{z}_{t}=\bm{A}\bm{z}_{t-1}+\bm{W}\Phi^{*}(\bm{z}_{t-1})+\bm{C}\bm{s}_{t}+\bm{h},$ (1)
with $\Phi^{*}(\bm{z}_{t})=\left[z_{1,t},\cdots,z_{M-P,t},\max(0,z_{M-P+1,t}),\cdots,\max(0,z_{M,t})\right]^{T}$, and $\bm{z}_{t}\in\mathbb{R}^{M}$ being the system’s latent state. To avoid redundant parameterization, we define $\bm{A}\in\mathbb{R}^{M\times M}$ as a diagonal matrix where entries corresponding to the linear (first $M-P$) units are set to zero, $\bm{A}=\mathrm{diag}(0,\ldots,0,a_{M-P+1},\ldots,a_{M}),$ so that linear self-connections are only assigned to the nonlinear units. The matrix $\bm{W}\in\mathbb{R}^{M\times M}$ captures interactions based on the partially nonlinear transformed state $\Phi^{*}(\bm{z})$, $\bm{C}\in\mathbb{R}^{M\times K}$ weighs the K-dimensional external inputs $\bm{s}_{t}$, and $\bm{h}\in\mathbb{R}^{M}$ is a bias term.
The AL-RNN partitions the state space into $2^{P}$ linear subregions separated by switching boundaries, within which the dynamics are linear and analytically tractable (see Fig. 9). Each linear subregion of the AL-RNN can be uniquely identified by a binary bitcode of length $P$, corresponding to the on/off (positive/negative) activation state of the $P$ ReLU units. Since the ReLU nonlinearity partitions each dimension at zero, this yields $2^{P}$ possible configurations of active units. By tracking which bitcodes occur during inference, we can quantify and visualize how many of these subregions are actually used by the network, offering a compact representation of its functional complexity.
This PWL structure offers a key advantage: it preserves linear dynamics within each subregion while still enabling the network to express nonlinear behavior through discrete transitions between linear subregions. This allows for efficient computation and clear interpretability. Transitions between subregions are explicitly detectable, and within each regime, the system’s dynamics can be analytically characterized in terms of fixed points and stability properties [15, 17] (see Appx. A.2). To retain interpretability, we constrain our models to a single AL-RNN layer throughout.
Training Details
To promote the reconstruction of both fast and slow time scales - and associated memory - in the latent space, we incorporate a regularization term based on Schmidt et al. [57] (see Appx. A.1.2 for details and performance comparisons). Across tasks, the AL-RNN is embedded in an encoder–decoder framework: raw sequences $\{\bm{x}_{t}\}_{t=1:T}$ are first mapped to latent inputs $\{\bm{s}_{t}\}_{t=1:T}$ via task-specific encoders. When the input modality has known structure (e.g., audio, vision, language), we apply appropriate nonlinear preprocessing (e.g., spectral representations, convolutions, or word embeddings, see Appx. A.1.3). For symbolic inputs, we directly use one-hot encodings. The AL-RNN processes these inputs, and task outputs are decoded from the latent states $\{\bm{z}_{t}\}_{t=1:T}$ or the final state $\bm{z}_{T}$ via a linear readout layer. While sophisticated nonlinear decoders could potentially offload complexity from the recurrent model, this consistent use of linear decoding isolates the contribution of recurrent dynamics to memory and task performance, ensuring that any long-range dependencies and nonlinear effects are captured within the recurrent structure itself.
Memory Mechanisms in RNNs
Sequential tasks require recurrent systems to store and manipulate information over time. RNNs implement these memory functions through distinct dynamical mechanisms that govern the evolution of hidden states, and associated computational mechanisms that determine how information is processed.
Linear RNNs exhibit a restricted but analytically tractable set of dynamics, governed by the spectral properties of the recurrence matrix [61]. Slow modes (eigenvalues near the unit circle) allow information to persist over time, while marginal stability supports long-lived activity in finely tuned cases. Oscillatory dynamics, from complex eigenpairs, enable rhythmic trajectories but require precise spectral alignment. The system is time-invariant and input-linear, lacking state-dependent transitions. These dynamics support basic memory operations such as leaky integration, graded retention, and temporal filtering [58, 54, 21]. Oscillatory modes - although fragile - can encode phase-sensitive information [27]. Linear RNNs cannot implement gating or context-sensitive routing, but can approximate input projection and superposition, enabling limited compositional processing.
Nonlinear RNNs support a broad repertoire of dynamical regimes, including multistability, where multiple attractors allow for distinct, stable internal states. Thresholded activations act as switches, enabling state-dependent transitions. More complex dynamics such as chaotic itinerancy and metastability generate structured, non-convergent trajectories. These dynamics support associative memory [29, 4], task and context-dependent switching [5, 67], and context-dependent routing, where identical inputs lead to different outcomes depending on internal state. Nonlinear gating allows selective integration or suppression of information [2]. Transient trajectories support sequential processing and structured recall [52, 64].
4 Experimental Results
Task Selection and Experimental Structure
To probe how recurrent nonlinearity shapes memory, we selected a set of tasks spanning those that can be solved by linear dynamics and those that require nonlinear computation, allowing us to highlight distinct linear and nonlinear memory mechanisms (Sect. 3) engaged by different computational demands.
4.1 Nonlinearity Reshapes Slow Modes in Temporal Integration
We begin with a class of tasks that require gradual accumulation of information over time to support a final decision. At each time step, the model receives a new input $\bm{x}_{t}$, and after a sequence of $T$ steps, it must predict a global output $\bm{y}$, such as a class label. We test this setup across diverse input modalities, including images, audio, and text. The setup mirrors the structure of drift diffusion processes in psychology [54], in which evidence is integrated over time until a decision threshold is reached. Analogously, the AL-RNN’s latent state $\bm{z}_{t}$ integrates incoming representations $\bm{s}_{t}$ over time, and a prediction $\hat{\bm{y}}$ (e.g., a class label) is made from the final state $\bm{z}_{T}$.
Sentiment Classification on IMDb
We first analyzed binary sentiment classification using the IMDb dataset [39]. Input sequences consist of tokenized movie reviews, which were standardized by truncation or padding to a sequence length of $128$ during training and testing. We mapped raw tokens $\bm{x}_{t}$ to latent inputs $\bm{s}_{t}$ by combining pretrained GloVe vectors [51] with a trainable nonlinear embedding layer. Fig. 1a illustrates the overall scheme. As depicted in Fig. 1b, varying the number of nonlinear units minimally affects performance, suggesting - as may be expected - a relatively simple linear temporal integration mechanism. During a negative review (Fig. 1d) large updates to the latent state occur almost exclusively when highly valenced keywords such as worst or terrible are encountered, confirming that memory integration is dominated by sparse lexical cues already emphasized by the input embedding. The distribution of nonlinear bitcodes is correspondingly degenerate (Fig. 1c): a handful of codes account for almost all examples.
Dynamically, temporal integration is driven by a dominant slow mode, reflected in a single eigenvalue near 1, with all others substantially smaller, indicating faster-decaying transients (Fig. 1e). This configuration enables the system to maintain and integrate information over time primarily along the slowest dynamical mode. Supporting this, the eigenvector associated with the slow mode is almost perfectly aligned (cosine similarity > 0.999) with the first PC of the latent trajectories, which alone accounts for $98\%$ of the total variance—indicating that the system’s activity is effectively constrained to a one-dimensional (sentiment-related) integration axis. Similar slow-mode or line attractor based integration mechanisms have been implicated in memory and evidence accumulation tasks across both artificial and biological systems [58, 40].
Figure 1: a: Schematic of the processing pipeline. b: Test accuracy as a function of nonlinear units, $P$. c: Bitcode frequencies revealing clustering of latent states. d: Trajectory through latent space showing sentiment-relevant keywords guiding classification. e: The dominant first PC clearly separates positive and negative reviews. The eigenvalue spectrum (right) has a maximum value close to one whose eigenvector aligns closely with the first PC.
Sequential MNIST (sMNIST)
Figure 2: a: Accuracy (y-axis) vs. number of nonlinear units (x-axis; mean $\pm$ std over 10 seeds). b: Final latent state projection onto the first 3 PCs show that nonlinearity partitions latent space by class.
In the sMNIST task, $28\times 28$ digit images are flattened into a sequential time series of length 784, with one pixel input per time step. Classification is based solely on the final latent state. To preserve sequential structure while enabling feature extraction, we apply a local 1D convolutional encoder before passing the sequence to the AL-RNN (Appx. A.1.3).
As for IMDb, we observe a similar slow accumulation of evidence along the first PCs, but now implementing a more complex, higher-dimensional data manifold (see Figs. 2&12). Accordingly, linear models already perform well on this task, with maximum eigenvalues slightly above 1 ($\lambda_{max}\approx 1.003\pm 0.003$, Fig. 12). However, as shown in Fig. 2a, increasing the number of nonlinear units still improves test accuracy up to an intermediate point, after which performance declines again for fully nonlinear models. Fig.2b shows final latent states projected onto the first three PCs, revealing how nonlinearity enhances class separation by partitioning the latent manifold, achieving up to $98.6\%$ test accuracy. Bitcode analysis provides a clear mechanism by which nonlinearity increases performance: each class aligns with its own set of closely neighboring bitcodes. For instance, for $P=50$, the average variation in the bitcode within each class was just $0.06\pm 0.02$ (see Fig. 13), meaning that on average $48$ out of $50$ PWL units had the same sign at the final state. This allows the AL-RNN to implement class-specific linear integration dynamics, resulting in well-separated clusters in latent space (Figs. 2b & 14). These results support the view that nonlinearity in this setting acts not as a core computational mechanism, but as a tool for improving expressivity. However, with full nonlinearity ($P=100$), this separating effect leads to more imbalanced variance distributions across classes (Fig. 14), degrading performance. We observed qualitatively identical behavior in an audio-based classification task (Speech Commands), reported in Appx. A.4.1.
4.2 Minimal Nonlinearity Stabilizes Transients For Structured Memory Recall
Next, to evaluate the AL-RNN’s ability to sustain stable internal representations over time, we tested it on the classic copy task (Fig. 3a). During the encoding phase, the network receives a random sequence of $N_{sym}$ distinct symbols, presented one per time step. This is followed by a delay of length $D$ time steps with no input, and then a “cue” signal indicates that the network should reproduce the original sequence over the next $N_{seq}$ time steps (the “recall phase”). Performance is measured as the number of sequences recalled correctly. This task is particularly interesting from a neuroscience perspective, as it parallels experimental paradigms used to study working memory in prefrontal cortex, where stable task-relevant information must be maintained over variable delay intervals despite complex and heterogeneous underlying dynamics [47, 53].
Figure 3: a: Top: Structure of the copy task and two example trajectories for $P=1$. b: Top: Activity of the PWL unit ($P=1$) for the same two input sequences. The latent activity follows a complex limit cycle primarily located in one linear subregion (PWL unit negative), which switches to the second subregion during decoding (PWL unit positive). Bottom: Autonomous activity of the AL-RNN in the absence of inputs encodes a 100-cycle, with its transient located only in one subregion (PWL unit negative). c: Symbol-wise recall accuracy (mean±$\pm$ std over 10 seeds) vs. number of PWL unit $P$. d: Histogram of binary “bitcodes” during recall for $P=10$, concentrated on a small subset out of $10242^{10}=$ 1024 possible bitcodes. e: Explained variance ratio of PCs of latent network activity (for $P=1$) indicates relatively high-dimensional, complex dynamics.
In this configuration $(N_{sym}=4,N_{seq}=8$ and $D=200)$, we first find that purely linear models perform above chance, suggesting they can partially solve the task. This aligns with theory: linear systems can support marginally stable oscillations when eigenvalues lie on the unit circle, enabling long recurrence times that could, in principle, encode sequence and timing (see Fig. 15). However, such solutions are highly fragile to noise and perturbations [61]. Minimal nonlinearity therefore greatly improves performance and robustness. A single nonlinear unit ($P=1$) often yields perfect recall, with performance declining for increased nonlinear units and being worst for fully nonlinear models (Fig. 3c). Crucially, we find that even fully nonlinear models often only use a small subset of available linear subregions (Fig. 3d, Fig. 16 for a full statistical analysis).
To understand the underlying computational mechanism, we analyzed a representative AL-RNN with $P=1$ that perfectly solves the task. We find that the PWL unit leverages two distinct linear subregions to separate encoding/decoding vs. storage dynamics. During the encoding and decoding phases, the network remains confined to the primary linear subregion optimized for integrating and recalling symbols (Fig. 3b, top). In contrast, during the delay phase, the network transitions to the second linear subregion, where it implements a complex transient cycle that is structured according to the encoded sequence.
The autonomous dynamics of the model establish a global k-cycle with a precise period of 100 time steps (exactly half the delay phase duration), as confirmed by a neutral maximum Lyapunov exponent ($\lambda_{max}\approx 0$, Fig. 3b, bottom). While the initial transient dynamics are constrained entirely to the lower subregion, the fully developed k-cycle spans both subregions, oscillating between them in a structured manner. This cycle is stabilized through the interplay of two distinct spectral properties: one with a contracting virtual fixed point ($\lambda_{max}\approx 0.992$) and the other around a diverging fixed point ($\lambda_{max}\approx 1.005$) which drives the initial transient dynamics. This structured alternation between contraction and divergence, causing periodic transients, allows the model to robustly maintain memory over long horizons without external input (Fig. 3b).
4.3 Nonlinearity Enables Gating in the Addition Task
Figure 4: a: Performance as a function of units $P$ (mean±$\pm$ std over 10 seeds). b: Latent trajectories in PWL space with $P=2$ nonlinear units. The second linear subregion is selectively activated only at the two masked time points ("input"), indicated by sharp transitions across quadrant boundaries. Outside these events, the network remains in a single linear regime, leading to smooth integration of the cumulative sum. c: Time course of masked inputs (red), network outputs (grey/black), and target values (blue) for two example trials, where the input occurs either early (light grey) or late (black) in the sequence. In both cases, the trajectory initially follows a linear drift prior to input, then transitions sharply to an elevated integration path with the same slope—offset to reach the target after 100 steps.
Next, we consider the addition problem [28], an established test of selective memory. The input consists of a continuous stream of random numbers, along with a sparse binary mask indicating two marked time points. The goal is to output the sum of the two marked times, ignoring all other inputs. Crucially, the mask is not directly fed into a separate gating mechanism, but the AL-RNN must internally learn to accumulate inputs only at the relevant times. This requires the network to integrate incoming values differently depending on the input—something linear RNNs are incapable of.
Our pipeline accurately identifies that, without nonlinearity, the network fails on this task (Fig. 4a). In contrast, with just one nonlinear unit ($P=1$) the AL-RNN achieves orders-of-magnitude lower error. Fig. 4b illustrates that nonlinearity enables the emergence of an internal gating mechanism, where the AL-RNN learns to selectively "open" and "close" an internal integration pathway conditional on the input mask. To sum the inputs, the network leverages a slow linear drift that accumulates over the full sequence length (Fig. 4c), combined with selectively adding the gated inputs. This mechanism cleanly separates the functional contributions of linear and nonlinear units and aligns with theoretically derived optimal PWL solutions for this task previously described in the literature [46, 57]. Interestingly, as before, even with increased nonlinearity (Fig. 17), the network preserves this strategy, distributing linear integration and gating across two distinct small sets of subregions. However, fully nonlinear models fail to capture this simplicity effectively: their expressivity leads them to "overfit" the integration process, fragmenting the linear accumulation mechanism (Fig. 17). This "over-approximation" of a fundamentally simple process introduces instability, resulting in degraded performance (Fig. 4a).
4.4 Nonlinearity Enables Internal Task Switching Through Context-Dependent Routing
Figure 5: a: Task schematic: Depending on the initial context cue (blue or red), the network must classify the sign of the integrated input sequence with a reversed decision rule. b: Test accuracy as a function of nonlinear units, $P$. c: Final latent states for both contexts with a linear decision boundary (dotted line, red and grey regions indicate classes). Right: Latent trajectories over time (black for Class 1, red for Class 2) cumulatively integrate evidence. The initial context determines the activation of the PWL unit, selecting the linear subspace: the PWL unit is positive in Context 1 and mostly negative in Context 2, inverting the temporal integration process to enable linear separability.
We next assess the network’s ability to switch between tasks using a context-dependent integration task (Fig. 5a), a design inspired by behavioral paradigms in neuroscience for studying flexible decision-making [42]. Each trial begins with a one-hot context cue, followed by a sequence of scalar inputs ($T_{seq}=100$). This task requires the network to switch between two internal decision policies based on a contextual cue. Purely linear AL-RNNs cannot solve this task, and plateau at $50\%$ accuracy (Fig. 5b), independent of how many linear units are provided. In contrast, the task is almost perfectly solvable with just one linear and one PWL unit, achieving up to $96\%$ accuracy.
We illustrate the central mechanism with one successfully trained model with one linear and one PWL unit. Here, the PWL unit primarily operates in two distinct subregions, based on the initial cue (Fig. 5c). In Context 1, the nonlinear unit essentially mimics linear integration, residing within a single subregion for the entire sequence, while the linear unit accumulates the input symmetrically. In Context 2, the nonlinear unit transitions to a different subregion, effectively inverting the sign of evidence accumulation within the linear unit, achieving the required task switching, leading to linear separability of the final states (Fig. 5c, right). This mechanism exemplifies context-dependent routing: the nonlinear unit acts as a switch, directing identical inputs through distinct internal linear dynamics depending on the initial context. See Fig. 18 for flow field and example trajectories.
4.5 Nonlinearity Enables Task Switching in a Stimulus Selection Task in Joint Task–Neural Modeling
Figure 6: a: Schematic of the contextual stimulus selection task used in the PFC-1 dataset [55]. In each trial, animals responded to auditory stimuli based on either the spatial origin (left/right) or the pitch (high/low), depending on the current task rule. b: Latent trajectories from a task-trained 2D model for two identical stimuli under different rules, illustrating context-dependent separation at decision time by the subregion boundary (black horizontal line). c: Task performance across models of increasing nonlinearity (P) and different latent dimensions (M) (left) and misclassification counts by stimulus and block (right). d: Mean firing rate MSE with and without inputs. e: Joint training performance when modeling both task and spike data: test accuracy (left), mean spike rate MSE (middle left), and coefficient of variation (CV) MSE (middle right) and mean spike correlation (right) as a function of the number of nonlinear units. f: Mean activation over time of a representative nonlinear unit that differentiates rules and modulates stimulus-driven activity conditionally.
To test our approach on empirical neuroscientific data, we evaluated it on neural population recordings from the CRCNS PFC-1 dataset [55], which captures single-unit activity in rodent auditory cortex during a flexible stimulus selection task. In this paradigm, animals are required to alternate between two rules: selecting the appropriate response based either on the spatial origin of the sound or on its pitch (Fig. 6a). The two stimulus dimensions are presented simultaneously but independently, with either a low or high warble (pitch cue) played on both sides, and a pure tone (side cue) played either on the left or right. Depending on the current rule, the animal must attend to either the pitch (low = "go", high = "no-go") or the side (left = "go", right = "no-go") of the auditory input. This setup creates contextual ambiguity: for example, an identical stimulus combination of a high-pitched warble and a tone on the left requires different responses ("go" under the spatial rule, "no-go" under the pitch rule). Hence the task studies linear and nonlinear neural computations in context-dependent behavior, with sparse, variable recordings posing a realistic challenge. Preprocessing steps and selection criteria are detailed in the Appx. A.3. To ensure a consistent task representation, we restricted training and evaluation to the $723$ trials in which the animal made the correct choice.
Task-training the AL-RNN
We first trained an AL-RNN to solve the task based purely on structured temporal input. Each trial mimicked one of the 723 empirically observed trials, beginning with a brief context cue indicating the active rule (spatial or pitch), presented for 100 ms (two time steps), followed by a one-second delay. Then, an auditory stimulus was presented (one-hot encoded for pitch and side), and the model was required to produce a binary decision (go/no-go), which was read out linearly from the latent state at the empirically observed decision time $\bm{z}_{t_{dec}}$ (see Appendix A.3). Despite the task’s contextual nature, we found that a 2D model was already often sufficient to linearly separate identical stimuli based on their preceding rule cue, and linear 3D models often solved the task perfectly. Systematic evaluation confirmed that increasing nonlinearity beyond this dimension did not yield substantial gains in task accuracy alone (Fig. 6c, left), and that errors were largely restricted to contextually ambiguous stimuli (e.g., Right–Low under the spatial rule; Fig. 6c, right).
Task Information Improves Spike Reconstruction
We next used the same AL-RNN to jointly model both the animal’s task behavior and the neural spike trains. That is, the network received identical inputs as in the task-only training—context cue, delay, and stimulus—and was required not only to produce the correct decision at the appropriate time, but also to reproduce the observed spiking activity across neurons from the same latent states of the AL-RNN (Eq. 1). To achieve this, we trained the AL-RNN within the Multimodal Teacher Forcing (MTF) framework [10], which allows direct training on non-Gaussian spike data via a Poisson decoder. The decoder is hierarchically conditioned on a 5-dimensional, trial-specific feature vector [11], enabling it to capture slow drift in spike statistics while preserving a shared task representation in the AL-RNN. See Appx.A.1.4 for training details and Fig.20 for an analysis of the learned feature vectors.
We found that including task information markedly improved spike reconstruction (Fig. 6d). Providing the model with the rule cue and stimulus enabled it to better predict the neural data, suggesting a strong alignment between latent decision dynamics and observed neural variability. Quantitatively, spike rates and coefficients of variation were well captured (see Fig. 19a). To evaluate the model’s temporal reconstruction, we correlated recorded and generated spike trains, yielding a mean correlation of $r=0.359\pm 0.002$. This approached the correlation between separate Poisson samples drawn from the same latent trajectory ($r=0.371\pm 0.057$), suggesting an upper bound imposed by the intrinsic variability of the Poisson process.
Nonlinearity Supports Integration Under Dual Constraint
In the joint setting, nonlinearity began to play a more significant role. As shown in Fig. 6e, increasing the number of nonlinear units improves both task accuracy and spike reconstruction ($r\approx-0.37,p=0.005$, with lower firing rate MSE and higher task accuracy being better). In contrast, linear models plateau around $87\%$ accuracy and exhibit an almost identical error pattern: they consistently respond “go” to all trials with left or low cues, regardless of the active rule, thus failing to resolve the contextual ambiguity (see Fig. 6c, right). In many such nonlinear models, we observed the emergence of single units that effectively gated the task rule. These units became active in one rule context and inactive in the other, with stimulus integration occurring only within the active subregion, providing an interpretable mechanism for context-dependent switching (Fig. 6f). Quantitatively, we examined the PWL subregion active at the time the input is first presented (after one second) and compared which subregions were used in blocks 2 and 4 (the two conditions where both spatial and pitch cues are present but different rules are active). In successful models ($>99\%$ accuracy), trials from these blocks were routed into almost entirely distinct subregions (on average, $97\%$ of trials fell into block-specific linear subregions), while in unsuccessful models this separation was much less pronounced (only $32\%$). Accuracy was strongly correlated with this separation score ($r=0.66$). This shows that successful models leverage nonlinearity to assign different task rules to distinct linear regimes, allowing them to react differently to identical stimuli depending on context. This mechanism is illustrated for a successful task-trained 2D model in Fig. 6b.
Our results underscore an interesting tension in jointly modeling task behavior and neural activity: while the task is linearly solvable in 2D when trained on the task alone, introducing the sparse, noisy and variable cortical spike data, imposes additional complexity that benefits from the additional expressivity provided by nonlinear models.
4.6 Nonlinear Decoding Enables Compositional Generalization
Finally, we investigate the SCAN benchmark [33], which requires to systematically recombine learned action primitives according to syntactic structures, tapping into executive functions such as manipulation and control of memory content. In this setting, a sequence of linguistic commands (e.g., run opposite left after walk right) must be mapped to a corresponding sequence of low-level actions (e.g., turn right, walk, turn left, turn left, run; Fig. 10). The SCAN grammar is composed of conjunctions (and) and temporal dependencies (after), requiring the model to either execute commands sequentially or invert their order during decoding [33].
We adopt a modular encoder–decoder architecture, assigning a separate AL-RNN to encode the commands and to decode the resulting latent state into action sequences, separately varying the number of nonlinear units $P$ in the encoder and decoder AL-RNNs (out of a total of 128 units each). As for the two previous tasks, purely linear models remain stuck at around $20-30\%$ sequence accuracy (Fig. 7a). Sparse nonlinearity ($P=1,4$) - while improving performance - remains insufficient to solve the task reliably. Performance peaks when both encoder and decoder include moderate nonlinearity ($P=32$), with near-perfect accuracy ($>99.9\%$), and then slightly deteriorates for fully nonlinear models ($P=128,\approx 96\%$).
In addition, we observe a clear asymmetry: a linear encoder can suffice, achieving up to $99\%$ accuracy when paired with a sufficiently nonlinear decoder ($P=64$), while a linear decoder gets stuck below $80\%$ regardless of the encoder. To better understand this behavior, we analyzed the latent structure produced by the encoder. Even when the encoder is purely linear, the final encoded states map onto a large number of distinct subregions in the decoder’s PWL space: for instance, in a model with $P_{\mathrm{enc}}=0$ and $P_{\mathrm{dec}}=64$, the 4,182 test commands map to 4,054 distinct decoder subregions. Thus, while the encoder itself remains linear, end-to-end training organizes the latent representations such that the decoder’s nonlinearity can be exploited effectively.
Second, the combination of a fully nonlinear encoder ($P=128$) and a moderately nonlinear decoder ($P=4$) can also achieve $>99\%$ test accuracy while enabling a clear syntactic interpretation of the latent units. We found that this behavior emerges from the interaction between the expressive encoder, which maps syntactic constructs onto the PWL latent space of the decoder, effectively leveraging its subregions for decoding (Fig. 7). Specifically, a compact subspace of four subregions, characterized by the first two PWL units remaining inactive, captures the majority of intial states corresponding to shorter action sequences (Fig. 7b). In contrast, more complex and lengthier command sequences are consistently mapped to a separate latent subspace, where nonlinear dynamics become more pronounced (Fig. 7c). Within this second subspace, we observe a clear syntactic segregation, with commands containing and and after being mapped to mutually exclusive linear subregions (Fig. 7d). Additionally, we find that during decoding, transitions between composite commands in an action sequence are synchronized with sign flips in the PWL units (Fig. 21).
Collectively, these findings suggest that while symbolic encoding in SCAN can remain largely linear, compositional generalization depends on the PWL structure of the decoder that tiles the latent space into modular, context-sensitive, and syntax-aware subregions. This decomposition reveals a pathway toward interpreting how abstract syntactic operations are organized in more complex sequence tasks.
[…]
5 Discussion
This work systematically investigates how nonlinearity supports memory and structured computation across a diverse range of sequence modeling tasks. Using AL-RNNs, we identify task-dependent thresholds where sparse nonlinearity dramatically enhances expressivity and memory capacity. Across tasks, simple linear mechanisms form the backbone of memory, while selective nonlinearity enables task and context-specific computations. The PWL structure of the AL-RNN makes these mechanisms transparent, directly linking switches in model dynamics to computation. Our findings challenge the assumption that greater nonlinearity always improves expressivity. We show that fully nonlinear models often perform worse than a combination of linear and nonlinear units. This may partially explain the efficiency of linear SSMs in tasks with long-range dependencies [24, 60, 48], and highlights the benefits of a targeted, sparse approach to nonlinearity in recurrent models.
The perspective that the structured separation of linear and nonlinear dynamics enables the extraction of computational mechanisms is also highly relevant for analyzing the latent structure of neural dynamics [16], which encodes similar mechanisms such as storage, gating, and symbolic manipulation [67, 42, 12]. Our analysis of rodent neural data supports this view: although the task is linearly solvable, jointly modeling behavior and spike activity benefits from nonlinearity to accommodate the variability inherent in empirical neural data. Our results demonstrate how structured models like the AL-RNN can uncover interpretable dynamics in noisy, real-world settings, while also highlighting that much of the complexity in neural data may actually stem from noise, drift, or unobserved inputs, not the task itself.
We focus primarily on memory in RNNs but also demonstrate that operations relevant to reasoning, such as conditional processing and context sensitivity, emerge naturally from the underlying PWL dynamics. Mechanisms like gating and symbolic recombination are enabled by specific nonlinear transitions, illustrating how RNNs can perform structured reasoning without explicit symbolic representations. The modular organization of latent space through PWL units, akin to cortical processing [20, 3], offers a structured pathway for mechanistic interpretability.
Limitations and Future Directions
Our study covers a structured but limited set of memory tasks. While the identified dynamics capture important memory strategies, many real-world tasks, particularly in language, may blend these mechanisms more intricately. Extending our analysis to more complex architectures, including structured SSMs like Mamba [22] or Transformer-based models [65], represents a promising path forward for bridging interpretability and SOTA sequence models.
Acknowledgements
This research was funded by the German Research Foundation (DFG) within the framework of the Collaborative Research Centre subprojects A06 and B08. We thank the Hector II Foundation for additional support.
Appendix A Appendix
A.1 Methodological Details
A.1.1 Training Details
Optimization and Hyperparameters
The primary focus of this study was to investigate how nonlinearity affects performance. To ensure that our observations were not confounded by latent dimensionality, we used relatively high-dimensional AL-RNN models, scanning over several values of the latent size $M$ in the range of 10 to 100 (up to $128$ for SCAN given the higher complexity of the task). This allowed us to operate in a regime where performance was not limited by the latent state capacity. The main hyperparameter of interest was the number of nonlinear units $P$, which was systematically varied across experiments to assess its impact on task performance and dynamical properties. Hyperparameter scans for $P$ are reported directly in the results figures of the main text.
For optimization, we employed standard settings consistent with the original AL-RNN implementation [9], including initialization strategies from the official repository. We initialized the self-connection weights in $\bm{A}$ with small random values to prevent premature convergence to unstable dynamics, while the weights $\bm{W}$, the bias term $\bm{h}$, the input matrix $\bm{C}$ and readout matrix $\bm{D}$ (mapping to logit scores for classification) were sampled from a Gaussian distribution with zero mean and a standard deviation of $0.01$. For optimization, we employed Adam [32] with a learning rate of $0.001$, which was further reduced during training by a cosine annealing learning rate scheduler. We used these settings for all datasets and did not perform fine-grained tuning.
During training, the model parameters were trained using standard backpropagation through time (BPTT). An independent validation set, comprising $10\%$ of the training data, was used for model selection. For datasets with pre-existing validation splits (e.g., sMNIST, Speech Commands, IMDb, SCAN), this was taken from the training set. For datasets without predefined splits, a validation set was artificially created. Other hyperparameters, such as batch size and learning rate scheduler, were set to conventional values that ensured stable convergence across all tasks.
Hardware Usage
All models presented in this study are relatively lightweight, using single-layer AL-RNNs with a maximum of 128 units. Due to this efficiency, all experiments could be conducted on a single CPU, with total training times for each model not exceeding 12 hours. Hence, the memory footprint remained well within the capacity of standard computing setups (less than $8$GB RAM). The modest computational requirements ensure that our approach is easily replicable on conventional hardware.
A.1.2 Manifold Attractor Regularization
To promote the reconstruction of both fast and slow time scales - and associated memory - in the latent space, we regularize a subset $M_{\mathrm{reg}}\leq M$ of the latent states, as suggested in Schmidt et al. [57], according to:
$\mathcal{L}_{\mathrm{reg}}=\tau\left(\sum_{i=1}^{M_{\mathrm{reg}}}(\tilde{A}_{ii}-1)^{2}+\sum_{i=1}^{M_{\mathrm{reg}}}\sum_{j\neq i}^{M}(W_{ij})^{2}+\sum_{i=1}^{M_{\mathrm{reg}}}h_{i}^{2}\right).$ (2)
Here, $\tilde{\bm{A}}$ represents the effective diagonal term for each unit: for linear units, the diagonal elements are derived directly from the $\bm{W}$ matrix, as $\bm{A}$ is zero in this case. For nonlinear units, the diagonal contribution is computed as the sum of their respective entries in $\bm{A}$ and $\bm{W}$. This ensures that the regularization correctly constrains self-connections to be near $1$, irrespective of the unit’s type. The regularization loss is added to the total loss with a scaling factor $\tau$ for each setting, which we scanned for $\tau\in\{0,0.1,0.5,1.0,5.0,10\}$ for the integration tasks (Fig. 8). We found a value of $0.1$ to overall lead to good performance, which we settled on it for all results presented in the paper.
By constraining the self-connections of certain units to be near 1 and suppressing their cross-connections, these sub-units function as near-perfect integrators, capable of encoding arbitrarily slow time scales. Dynamically, this regularization encourages the formation of a continuous manifold of marginally stable fixed points, supporting long-term integration without requiring architectural modifications such as gating. It enables efficient memory retention even in simple RNN architectures, and has been shown to outperform LSTMs and other RNN architectures on long-range tasks [57]. As recommended by Schmidt et al. [57], we applied the regularization to half of the latent units $M$, hence primarily targeting the linear subspace, depending on the number of $P$ PWL units.
Figure 8: Manifold Attractor Regularization (MAR, [57]) leads to a strong performance boost on sMNIST and Speech Commands Task, while not being overly sensitive to the exact value of the regluarization strength $\tau$.
A.1.3 Encoders for Text, Audio, and Images
To incorporate domain-specific inductive biases, we used custom encoders for each modality, optimized for capturing the relevant temporal and structural patterns inherent in the data.
IMDB
For the IMDB sentiment classification task, we use pre-trained $100D$ GloVe embeddings [51] to initialize word vectors, followed by a multi-layer perceptron (MLP) as a nonlinear encoder. This additional nonlinear mapping allows the model to capture high-level semantic abstractions from raw word embeddings, which is crucial for sentiment classification [31, 69]. Without this structured transformation, the model struggles to resolve complex syntactic and semantic relationships in the text.
sMNIST
For the sMNIST task, we apply a series of stacked local 1D convolutional layers to the pixel sequences. These 1D convolutions respect the sequential nature of the input while extracting local stroke-based features that are beneficial for digit recognition.
Speech Commands
For the Speech Commands dataset, we transformed the raw audio signals into Mel Frequency Cepstral Coefficients (MFCCs), a widely used feature representation for speech processing [14]. MFCCs provide a compact and noise-robust representation, emphasizing the perceptually relevant aspects of speech. This transformation also downsamples the sequence length based on the MFCC hyperparameters. In our setup, we select parameters that preserve moderate sequence lengths comparable to sMNIST $16k$ time steps. The transformed MFCC features are then passed through a series of stacked local 1D convolutional layers to extract temporal patterns. To ensure consistency across samples, the MFCC features are normalized using global mean and standard deviation, computed only over the training data.
A.1.4 Training on Spike Data
To train on the neural spike data in Sect. 4.5 and Appx. A.3, we embedded the AL-RNN within the Multimodal Teacher Forcing (MTF) framework [10], which enables training directly on non-Gaussian observations through likelihood-based optimization. Specifically, we employed a 1D temporal convolutional encoder to map observed spike trains to the latent space, followed by a Poisson decoder defined as
$\log p(\bm{c}_{t}\mid\mathbf{z}_{t})=\sum_{i}\left(c_{ti}\log\lambda_{ti}-\lambda_{ti}-\log(c_{ti}!)\right)\quad\text{with}\quad\lambda_{t}=\exp(\bm{B}\bm{z}_{t}+\bm{b}),$
where $\bm{z}_{t}$ is the latent state of the AL-RNN and $\lambda_{t}$ the predicted spike rate and $\bm{c}_{t}$ the observed counts. The MTF framework incorporates a latent consistency loss that penalizes the distance between inferred and predicted latent trajectories by
$\mathcal{L}_{\text{consistency}}=\sum_{t}\|\bm{z}_{t}^{(\text{enc})}-\bm{z}_{t}^{(\text{gen})}\|^{2},$
where $\bm{z}_{t}^{(\text{enc})}$ is the latent trajectory inferred from the encoder and $\bm{z}_{t}^{(\text{gen})}$ the trajectory generated by the AL-RNN. Given the short duration of the spike train trials (30–40 time steps) and their high noise levels, we found that training without teacher forcing but relying solely on the consistency loss yielded the best results, likely because teacher forcing caused overfitting to individual spike events.
To account for drift in trial-specific spike statistics while preserving a shared task representation, we hierarchically parameterized the decoder following the approach in [11], but applied this structure exclusively to the decoder parameters. Specifically, each trial was associated with a five-dimensional feature vector $\bm{l}^{(j)}$ (with its dimension determined by PCA, Fig.20) that generated trial-specific decoder weights and biases via learned linear projections. The resulting hierarchical Poisson decoder was defined as
$\log p(\bm{c}_{t}\mid\bm{z}_{t},\bm{l}^{(j)})=\sum_{i}\left(c_{ti}\log\lambda_{ti}^{(j)}-\lambda_{ti}^{(j)}-\log(c_{ti}!)\right),\quad\text{with}\quad\bm{\lambda}_{t}^{(j)}=\mathrm{softplus}(\bm{B}^{(j)}\bm{z}_{t}+\bm{b}^{(j)}),$
where $\bm{B}^{(j)}$ and $\bm{b}^{(j)}$ are decoder parameters generated as linear projections from $\bm{l}^{(j)}$.
This design choice ensures that the AL-RNN learns a consistent dynamic structure across trials. Hierarchizing the RNN itself would allow it to absorb task variation into the dynamics. By restricting flexibility to the output mapping, the AL-RNN remains responsible for solving the core disambiguation problem imposed by the context-dependent task, while the decoder accounts for slow drift or baseline shifts in firing statistics. Although the decoder is not tasked with solving the decision problem, the learned feature vectors nonetheless capture systematic variation linked to task conditions. Specifically, PCA of the trial-specific feature vectors revealed that the first three components were sufficient to linearly separate go vs. no-go trials with $87\%$ accuracy (Fig. 20). A similar level of separability ($\sim 85\%$) was observed when applying PCA directly to the trial-averaged spike data, indicating that the central moments of the firing statistics already encode task-relevant differences. The hierarchical decoder leverages this structure to improve spike reconstruction, but without access to the input structure required to disambiguate the task itself.
Across all trials, we observe a total of 783,840 time points. For the results presented, we use an AL-RNN with $M=10$ latent dimensions, yielding 220 trainable parameters for the recurrent model. The hierarchical decoder includes $N_{\text{feat}}=5$ learnable features per trial, which, combined with the shared linear projections, results in a total of 5,375 trainable parameters. Given the large number of observed data points relative to the model’s capacity, this setup remains well within a regime where overfitting is unlikely.
A.2 Analysis of Trained Models
Bitcode Analysis
To quantify the distribution of latent states across the linear subregions, we extract the last $P$ components of each latent state trajectory and represent their activation patterns as bitcodes, where each component is assigned a binary value based on its sign (1 if positive, 0 otherwise). Formally, the bitcode for a latent state vector $\bm{z}_{t}\in\mathbb{R}^{P}$ at time step $t$ is defined as:
$b_{t}=\sum_{i=1}^{P}\mathbb{I}[z_{t,i}>0]\cdot 2^{P-i},$ (3)
where $\mathbb{I}[\cdot]$ is the indicator function. The empirical distribution of these bitcodes across a set of latent states is then given by:
$p(b)=\frac{n(b)}{\sum_{b^{\prime}\in B}n(b^{\prime})},$ (4)
where $n(b)$ is the count of bitcode $b$ and $B$ is the set of all observed bitcodes.
Figure 9: Illustration of the AL-RNN and bitcode assignment. The example displays a 6-dimensional AL-RNN with four linear and two PWL units (left). The PWL units partition the state space into four linear subregions (right). Bitcodes encode positive $(1)$ and negative $(0)$ activation values for these units.
Stability Analysis
In each linear subregion of the AL-RNN, its fixed points and their stability can be analytically computed. For a given bitcode, we first compute the masked recurrent weight matrix:
$\bm{W}_{\mathrm{masked}}=\bm{W}\odot\bm{M}(\text{bitcode}),$ (5)
where $\bm{M}(\text{bitcode})$ applies masking to the connectivity matrix $W$, effectively zeroing out the connections of inactive nonlinear units.
Fixed points are then computed by solving the linear system:
$(\bm{A}+\bm{W}_{\mathrm{masked}}-\bm{I})\bm{z}^{*}=-\bm{h}$ (6)
using a standard linear solver (implemented with numpy.linalg.solve). If the matrix is singular, no fixed point is considered valid.
To assess the stability of each fixed point, we compute the eigenvalues of the Jacobian matrix:
$\bm{J}=\bm{A}+\bm{W}_{\mathrm{masked}}.$ (7)
The fixed point is classified as stable if all eigenvalues $\lambda$ satisfy $|\lambda_{i}|<1$, which were computed via the eigenvalue decomposition using numpy.linalg.eigvals.
Maximum Lyapunov Exponent
The maximum Lyapunov exponent, $\lambda_{\max}$, characterizes the average rate of divergence of infinitesimally close trajectories in a dynamical system. It is formally defined as:
$\lambda_{\max}=\lim_{T\to\infty}\frac{1}{T}\log\left\|\prod_{r=0}^{T-2}\bm{J}_{T-r}\right\|$ (8)
[…]
Variance Analysis
To quantify the spatial distribution of class manifolds in the AL-RNN latent space for the sMNIST task, we computed four metrics based on the final latent states of each digit class (Fig. 14). The final latent states were first normalized dimension-wise to account for scaling differences and then projected onto their PCs. The number of components was dynamically selected to capture at least $80\%$ of the total variance. We first evaluated the Coefficient of Variation (CV), which is defined as the ratio of the standard deviation to the mean of the class-specific variances $\bm{v}$: $\mathrm{CV}=\frac{\sigma(\bm{v})}{\mu(\bm{v})}$. The CV provides a measure of relative dispersion, indicating how spread out the class-specific variance is in relation to its average. We further computed the Gini coefficient, which captures inequality in the distribution of variance across class manifolds, given by $G=\frac{\sum_{i=1}^{N}\sum_{j=1}^{N}|v_{i}-v_{j}|}{2N^{2}\cdot\mu(\bm{v})}$. The Gini coefficient specifically quantifies how unevenly the variance is distributed among the different classes. Third, we calculated the Max-Min Ratio, defined as the ratio of the maximum to the minimum class variance, defined as $R_{\max/\min}=\frac{\max(\bm{v})}{\min(\bm{v})}$. This ratio highlights the degree of disparity between the most and least represented class manifolds, providing a direct measure of distribution extremes. Finally, we measured the Shannon Entropy of the variance distribution: $H(\bm{v})=-\sum_{i=1}^{N}p_{i}\log(p_{i})$, where $p_{i}$ represents the normalized variance for each class. Lower entropy values imply that variance is concentrated within a limited number of classes, while higher entropy suggests a more even spread across all class manifolds.
Flow Field Approximation
To approximate the continuous latent dynamics of the AL-RNN, we computed the flow field over a two-dimensional grid of points in the latent space. At each grid point, we applied the AL-RNN’s step (Eq. 1) to estimate the local velocity vectors. This provides an approximation of the underlying continuous flow. For Fig. 18, we used a grid density of $30\times 30$ points.
A.3 Datasets
IMDb
The IMDb dataset [39] consists of 50,000 movie reviews, evenly split between training and test sets, with sentiment labels (positive or negative) for binary classification. It is available on Kaggle at https://www.kaggle.com/datasets/lakshmi25npathi/imdb-dataset-of-50k-movie-reviews. Each review was tokenized and converted to word vectors using pre-trained GloVe embeddings [51]. The sequences are truncated or padded to a fixed length of 128 tokens during training and testing. We did not apply any further text preprocessing beyond tokenization, as the GloVe representations handle most standard normalization. Training and evaluation follow the standard 25,000/25,000 train/test split.
sMNIST
The sequential MNIST (sMNIST) dataset transforms the standard MNIST [34] digit images into sequences by flattening each $28\times 28$ image into a 1D sequence of length 784. We used the dataset as provided in torchvision.datasets [41]. The pixel intensity values are normalized to the range $[0,1]$ before being passed to the AL-RNN. The dataset is split into $60,000$ training sequences and $10,000$ test sequences, following the standard MNIST partitioning.
Speech Commands
The Speech Commands dataset [68] consists of 1-second audio recordings with a sampling rate of 16kHz of spoken words from a fixed vocabulary of commands: ["down", "go", "left", "no", "off", "on", "right", "stop", "up", "yes"]. We used the version of the dataset provided by TensorFlow Datasets [1], which includes standardized preprocessing, splitting, and metadata.
Copy Task
The Copy Task is a synthetic memory benchmark where the model is required to store and reproduce a random sequence of discrete symbols after a delay period. Each input sequence consists of a random sequence of symbols represented in one-hot encoding, followed by a blank segment of the same length, and finally a cue signal indicating when to begin recall. The target output is an exact reproduction of the original symbol sequence during the recall phase. Sequence lengths are fixed during training. The training set comprised $1000$ sequences, while the test set contained $200$ sequences. Given that the task is defined over $4$ symbols and $8$ time steps, the total number of unique sequences is $5364^{8}=65,$ 536. This represents a substantial combinatorial space, ensuring that the relatively small training set could not fully cover the distribution of possible sequences, thereby preventing the model from simply memorizing the training examples. Reported test accuracies reflect the percentage of correctly predicted symbols in the test sequences after the delay period.
Addition Problem
In the Addition Problem, popularized in [28], the goal is to sum the values at the marked positions and output the result at the end of the sequence. This task tests the model’s ability to integrate sparse, context-dependent information across long temporal spans. The two randomly chosen indices indicating the values to be summed are presented in the first half of the sequence. The AL-RNN is trained to produce the sum of the two digits at the final time step. The training set comprised $2000$ examples, while the test set contained $200$ examples. Reported test errors quantify the mean-squared error between the predicted and the correct sum.
Contextual Multistability
The Contextual Multistability task is a custom-designed integration task inspired by neuroscience paradigms of context-dependent decision-making [42]. Each trial begins with a one-hot context cue, indicating which integration policy the model should adopt. A sequence of noisy evidence is presented across several time steps, and the model is required to integrate this evidence according to the initial context. At the final time step, a recall cue triggers the model to produce a decision based on the integrated evidence. If the initial context is inverted, the decision boundary is reversed. The AL-RNN is trained to predict the correct label based on its final latent state. We trained on $1000$ example sequences, and tested on $200$ sequences.
Prefrontal Cortex Task Recordings
The CRCNS PFC-1 dataset [55], includes single-unit recordings from medial prefrontal cortex and auditory cortex of rats performing a context-dependent stimulus selection task. We focused on the longest available session (Day 4) from the first rat (CR12B), comprising approximately 900 trials. For this rat, only auditory cortex (AC) neurons were recorded. Each trial involves auditory stimulus comprising a pitch component (high or low warble, presented bilaterally) and a spatial cue (broadband noise from left or right). The correct response—"go" or "no-go"—depends on the current task rule: either localization (respond based on side) or pitch discrimination (respond based on pitch). The task is structured in repeating blocks. The task was structured into repeating blocks of four trial types that defined the task stage.
  1. 1.
    Block 1: spatial rule with spatial-only stimuli (left vs. right broadband noise),
  2. 2.
    Block 2: spatial rule with compound stimuli (both spatial and pitch cues)
  3. 3.
    Block 3: pitch rule with pitch-only stimuli (low vs. high warble, bilaterally presented)
  4. 4.
    Block 4: pitch rule with compound stimuli
We aligned each trial from one second before stimulus onset to the decision point (center-port withdrawal for go trials). No-go trials lacked an explicit decision timestamp, so we truncated them to match the duration distribution of go trials, avoiding length-based confounds between trial types. Spike times were binned at 50ms resolution, resulting in spike count vectors for each neuron over the trial window.
External inputs were encoded as one-hot vectors: two dimensions for the spatial cue (left/right), two for pitch (low/high), and four for the current block (task stage). To mimic the contextual cue structure used in memory tasks, task stage information was only provided during the first 100 ms of the trial, reflecting the fact that rats had learned the current rule through block structure and early trial cues. Only trials with correct responses were retained to ensure that task-trained and spike responses were consistent with each other.
SCAN Task
The SCAN dataset [33] is a synthetic benchmark designed to evaluate systematic compositional generalization in sequence-to-sequence learning. In SCAN, natural language commands are translated into sequences of low-level actions in a navigation environment. For example, the command "jump twice" is mapped to the action sequence "JUMP JUMP", while a more complex command like "walk around right" is interpreted as "RTURN WALK RTURN WALK RTURN WALK RTURN WALK". The mapping is deterministic and unambiguous, with each command generated by a phrase-structure grammar that defines valid language-action mappings. The commands are constructed from a small set of primitives:
  • Primitive Actions: walk, run, jump, look.
  • Directional Modifiers: left, right, opposite.
  • Frequency Modifiers: twice, thrice.
  • Compositional Operators: and, after.
We specifically use the Simple Split of the SCAN dataset, where $80\%$ of all command-action pairs are randomly assigned to the training set, while the remaining $20\%$ are reserved for evaluation. The primary challenge is to generalize compositional patterns from seen instructions to novel ones. For instance, if the model learns "walk" and "jump", along with the modifier "twice", it should generalize to "jump twice" even if that specific combination was not observed during training. We follow the canonical implementation of SCAN for preprocessing and evaluation, where commands are tokenized into one-hot vectors. During training, input sequences are encoded with the encoder AL-RNN and predicted using the decoder AL-RNN, which runs freely from the initial state provided by the encoder AL-RNN. Reported test accuracies quantify the amount of correctly predicted test sequences.
Figure 10: Outline of the the encoder-decoder architecture used for solving the SCAN task.
A.4 Further Results
A.4.1 Speech Commands
To verify that the trends observed in sMNIST generalize across modalities, we evaluated the AL-RNN on the Speech Commands dataset [68], a ten-class audio classification task using spoken commands sampled at $16$ kHz.
We used the same evaluation setup as for sMNIST, with classification based on the final latent state after processing the input sequence. As in the visual domain, we observed that increasing nonlinearity improved classification accuracy up to an intermediate number of nonlinear units, after which performance declined for fully nonlinear models (Fig. 11a). Latent states from models with moderate nonlinearity again exhibited clear class-specific clustering in PC space (Fig. 11b), and bitcode alignment within classes remained high (Fig. 13)—closely mirroring the gating behavior observed in sMNIST (Sect. 4.1). These results confirm that the latent integration and partitioning mechanisms learned by the AL-RNN generalize robustly across visual and auditory modalities.
Figure 11: Same figure as Fig. 2 but for the Speech Commands task. a: Accuracy plotted against number of nonlinear units. Increasing nonlinearity leads to moderate increases, while performance deteriorates for fully nonlinear models. b: Final latent states projected onto the first 3 PCs highlight how nonlinearity partitions the latent space according to class labels.
Figure 12: a: Mean latent trajectories of the first PC for each digit class in the sMNIST task. The trajectories reveal a slow accumulation of evidence along the primary axis, similar to sentiment classification (Fig. 1). Interestingly, distinct branching patterns are visible: the digit ’6’ diverges first, reflecting its characteristic structure in the initial image segments. Digits ’8’ and ’9’ share overlapping paths until around the midway point, aligning with their visual similarity in the upper halves of the images. b: Explained variance ratio of the first $15$ PCs indicates a more complex structure compared to the $2D$ sentiment attractor in Fig. 1.
Figure 13: Analysis of class-specific bitcode activation for the sMNIST task (left) and Speech Commands (right). Left subpanel: Average bitcode activation probability per class for the final states $\bm{z}_{T}$ for two representative models of $P=50$ (top) and $P=100$ (bottom). High and low probabilities indicate that bitcode activations are consistent within a class. Right subpanel: Average within-class deviation from these predominant activation patterns (rows in the left figure). Notably, comparing $P=50$ (bottom) to the fully nonlinear model $P=100$ (top), the within-class variability of bitcode activations is reduced. This reduction in variation implies more stable and distinct bitcode representations, enhancing the robustness and reliability of the model’s classification outcomes.
Figure 14: Analysis of class manifolds for the sMNIST task. (a) Kernel Density Estimates (KDE) of key metrics across different nonlinear configurations ($P=0,50,100$). For the fully linear model ($P=0$), class variance is tightly packed and uniformly distributed, as indicated by low Gini coefficients and Max-Min ratios. Introducing moderate nonlinearity ($P=50$) allows the AL-RNN to strike an effective balance, with a CV close to $1$, signifying well-proportioned variance across class manifolds. This results in distinct spatial separability, as evidenced by lower Gini coefficients and more balanced Max-Min ratios. In contrast, the fully nonlinear model ($P=100$) shows larger Gini coefficients and very wide Max-Min ratios, indicating a much more uneven distribution of variance. This suggests that certain class manifolds (e.g., digit 5) are overly compressed while others are disproportionately spread out, as evidenced in (b): Visualization of the final latent states projected onto the first three PCs for digits 3 (orange) and 5 (green), which represent the most frequent class confusion. At $P=50$, class manifolds are spatially distinct and well-separated, indicating effective utilization of latent space. In contrast, for $P=100$, these manifolds become elongated and entangled, with digit 5 particularly over-compressed. This spatial entanglement directly correlates with the misclassification patterns observed, reflecting an inefficient partitioning of latent space.
Figure 15: Top: First PC of latent activity during the copy task for the best-performing linear model ($65\%$ test accuracy). Bottom: First 3 PCs of freely generated activity.
Figure 16: Distribution of linear subregions during decoding stage for the copy task. (a) The Gini coefficient of bitcode distributions as a function of $P$, reflecting the imbalance in subregion usage. As $P$ increases, the Gini coefficient rapidly rises, indicating that only a limited subset of regions is dominantly occupied. (b) The effective number of linear regions occupied by the AL-RNN, plotted alongside the theoretical maximum (dashed gray line) computed as $\min(2^{P},\text{total samples})$. Despite the potential to explore exponentially many regions, the network only populates a small fraction. (c) Cumulative mass plot for the bitcode distributions sorted by probability, visualized across different values of $P$. Larger $P$ results in most of the probability mass being captured by a small subset of regions.
Figure 17: Distribution of bitcode subregions in the addition task for different levels of nonlinearity $(P)$ for the best-performing models. Across all levels of nonlinearity, we observe a characteristic structure: a small subset of dominant subregions encode the linear integration dynamics, while sparse gating is implemented by a few low-frequency subregions (red). The relative ratio of subregion activity for masked time points is color-coded, and indicates a clear separation of subregions visited during linear integration (blue) and during gating (red), with no region implementing both tasks (see colorbar). While in the more nonlinear settings ($P=10,50$), the linear integration dynamics are distributed over multiple regions, these regions are close neighbors. This could introduce instability compared to the more compact representations seen with sparse nonlinearity, potentially explaining the performance decrease observed with fully nonlinear models. Note that for $P=10\&50$, the y-axis was scaled logarithmically for better readability.
Figure 18: Flow field and latent trajectories in the linear and nonlinear subspaces for both contexts (Context 1: blue, Context 2: red) in the contextual integration task. The left panel depicts the cumulative sum of evidence for two example sequences, with context inverting the decision rule. The right panel shows the vector field, where thick lines represent paths taken when no input is provided except the initial context cue, gradually drifting towards the decision boundary in the respective context. Thinner lines illustrate the two example trajectories under continuous input (corresponding to the cumulative sum of evidence shown in the left panel). In these cases, evidence integration drives the trajectories along context-specific pathways—positive accumulation in Context 1 and inverted accumulation in Context 2. An unstable fixed point (green dot) near the origin demarcates the boundary between the two decision regions.
Figure 19: a Predicted vs. true mean firing rate (top) and coefficient of variation (CV; bottom) across all neurons for an example model trained with task inputs ($M=10,P=3$). The model accurately captures both first- and second-order spike statistics. b Example raster plot comparing true and reconstructed spike trains for a single trial for the same model ($M=10,P=3$). c Example PCA-projected latent trajectories for the four ambiguous input conditions from Fig. 6, showing clear separation based on context.
Figure 20: a: Explained variance of the five PCs of the learned 5D trial-specific feature vectors, used to generate trial-specific decoder parameters. b: Predicted vs. true trial-averaged firing rates for all neurons. c: PCA projection of the feature vectors into the first three PCs, colored by Go (blue) and No-Go (red) labels. Trial-level task information is partially encoded in the decoder feature space; a linear classifier trained on the first three components achieves $86\%$ accuracy.
Figure 21: Latent Dynamics of PWL Units during composite actions in the SCAN task. a: Visualization of an example SCAN command walk around right thrice after jump around left thrice. Top: Executed actions over time, illustrating the switching behavior between distinct subtasks (look, turn left, walk). Middle: Activation of PWL Unit 1 throughout the sequence, highlighting a characteristic sign flip from negative to positive precisely before the transition point which triggers the switch to the next subtask (shaded background region). Bottom: Bitcode representation for the four PWL units, where black denotes negative activation and white represents positive activation. Notably, PWL Unit 1 consistently flips to positive at the switch, marking the boundary between subtasks. b: Statistical analysis of PWL unit activations around switch points between composite subtasks, averaged across all SCAN sequences. The x-axis represents time steps relative to the switch point ($t=0$). Bars indicate the average probability of positive activation for each PWL unit (mean $\pm$ sem). The dashed red line is the global baseline activation for each unit. All PWL units showcase characteristic shifts in their sign around switch points: Unit 1 initiates switches with a brief peak to positive activity, Unit 2 transitions from predominantly negative to predominantly positive activation (and vice versa for Unit 4), and PWL Unit 3 synchronizes its oscillations tightly with subtask boundaries.